# Blockchain Address Poisoning

### v1

@@ 8 Mitigations
Protocol-level mitigations: One mitigation is to map human-readable strings to each address, like DNS for IP addresses. Some entities already have adopted domain names on wallet addresses through the Ethereum Domain Name System (ENS). Human-readable addresses would reduce address poisoning as well as accidental transfers (Appendix C) and enhance usability. ENS, however, fosters other security risks such as typosquatting, name hijacking (i.e., re-using the expired domain names) [44], or name collisions [14].
@@
Contract-level mitigations: One solution is to modify the major stablecoin contracts and ask permission from the token sender even when the amount is zero. Function modification requires each ERC-20 token issuer to upgrade its smart contract despite limited incentives for developers. This solution only mediates zero-value transfers but not tiny and counterfeit token transfers. In addition, token owners can quickly blacklist the attacker’s address. For the largest payoff transfer we observe, the Tether team (USDT) blocked the attacker’s address within an hour, which prevented the attacker from moving the retrieved funds to another address. [^16] Transaction blocking $L$: 0x2675aaf5db84f9f966c431ff82f2173d3251020b11b4939b4961ebedf92c78dd.
@@ 9 Related work
We highlight work particularly related to our discussion, with a focus on attack scale. Gao et al. [8] identified 2,117 counterfeit tokens by applying keyword matching to the 100 most popular tokens. The authors investigated two techniques involving counterfeit tokens (airdrop and arbitrage), leading to total financial losses exceeding 17M USD over 7,000 victims. Xia et al. [43] identified around 10,000 counterfeit tokens on Uniswap (50% of all Uniswap tokens) through ground truth labels provided by the platform, address clustering, and machine learning-based techniques. They estimated these counterfeit tokens collectively led to financial losses of 16M USD from approximately 40,000 victims. He et al. [12] introduced TxPhishScope, a tool designed to detect blockchain phishing websites. The authors identified 26,333 phishing websites with 3,486 phishing blockchain accounts. Other studies on free giveaway scams reported 24M–69.9M USD losses [19], including 872K USD losses from Twitter [17]. Based on these other works, blockchain address poisoning appears to be one of the largest (in terms of aggregate losses) and most wildly targeted criminal activity taking place on-chain.
Closest to our efforts, Ye et al. [46] examined three types of blockchain scams, one of which is address poisoning on Ethereum with zero-value transfers. In parallel, and contemporary to our work, Guan and Li [10] also studied address poisoning on Ethereum and discuss all three poisoning transfers. We provide a detailed comparison with our method and results in §4.1, §5.2 and Appendix A. The Chainalysis’s blog post [34] conducted the case study (WBTC 68M loss) and analyzed its related addresses (victims, lookalike addresses).
@@ 11 Ethical considerations
The concept of blockchain address poisoning is already the subject of some online discussions, and, as this paper shows, this issue is currently extensively exploited in the field. As such, wallet operators and (some) users are most likely already aware of at least the theoretical possibility of the attack. What our work does, though, is to demonstrate its practical severity in economic terms – and we hope it will be a call to arms to act sooner rather than later.
While attackers could potentially misuse some of our findings to improve their strategies, besides raising community awareness, the main contribution of our paper is to extensively describe the attack mechanisms. We also introduce possible protection strategies. Our proposed mitigations can hopefully inspire wallet operators. Likewise, we hope that our work can help users recognize when they are in targeted populations, and accordingly, take precautionary measures.
Finally, Appendix G, specifically Table 9, includes full transaction hashes, which could potentially be used to re-identify the wallet addresses of the potential victims (whom we do not directly list in the table). We decided to present the table as is, as obfuscating part of the transaction hash would impede reproducibility or external validation of our findings, for little to no privacy benefit: users of Ethereum (or BSC) are relying on a public blockchain, in which all records are publicly available. Furthermore, wallet addresses are not, on their own, considered personally identificable information, as 1) there is no one-to-one mapping to specific individuals, as a given person may hold several wallets, and a single wallet may be shared by multiple people (e.g., belonging to the same organization), and 2) a wallet address does not, in itself, contain information directly mapping to a real-world identity.
@@ Acknowledgments
Appendix A Comparison with Guan and Li’s
This section compares our detection performance to the most recent work from Guan and Li [10]. To align with their observation, we only include our measurement results from Nov. 2022 to Feb. 2024 on Ethereum. Table 6 summarizes the results.
First, the total amount of payoff transfers is similar, corroborating the validity of both approaches. Second, both approaches estimate similar attack costs in ETH, but not in USD. Guan and Li use a fixed ETH to USD conversion rate (around 3200 USD) whereas we query the price at the time of attack. The exchange rate, however, fluctuated between 1000 USD and 2000 USD for the majority of the observation period, leading to the attack costs being overestimated by over 10 million USD, which results in a difference of a 300% on the return-on-investment (RoI) for the attacker. Third, we detect more than 2.2 million counterfeit token transfers. Guan and Li do not report tokens with symbols that do not exactly match USDT and USDC. For example, out of 5,838 counterfeit tokens, only 1,922 use the exact same token symbols whereas at least 740 tokens use the different encodings of USDT or USDC (e.g., using a Cyrillic “S” and/or “T” in “USDT”). Likewise, attackers can also avoid their detection if they change the attack characteristics, e.g., launching zero-value transfers without attack contracts. Our results suggest filtering the transfers based on address similarity first would be more robust to attacker manipulation (less false negatives), than filtering by attack characteristics. On the other hand, compared to Guan and Li [10], our technique misses attacks where less than 7 digits match the address $L$.
In addition to the methodological differences, our work has several distinctions. First, we have better coverage in measurement (two years, two chains). We find 270M attack attempts (compared to 14M), highlighting significant chain waste and user experience degradation.
@@ Appendix A Comparison with Guan and Li’s
Payoff Cost ETH ROI Tiny Zero-value Counterfeit CT transfer (USD) token Guan and Li [10] 76.79M 7918.2 (25.5M) 2.2 277.3K 7.85M 6.64M 2,333 Ours 76.81M 8202 (14.9M) 5.1 283.9K 7.05M 8.87M 5,838
Second, although Guan and Li also perform attacker clustering, they appear to have overestimated group sizes, due to erroneous merges from copying bots, as we show in §5.3. We carefully remove bots and extrapolate a few large attack entities.
Last but not least, most of our analyses are novel. We analyze victim characteristics, success conditions, group competitions, cross-chain attacks, and modeling of address generation for the first time. We provide deeper insights on how attackers perform attacks and what capabilities they have.
Appendix B Birthday Paradox
@@
Here, we examine whether our system should exclude a set of victims that present a collision probability greater than a given threshold $a$, i.e., $p\geq a$. We denote this set of victims by $\mathcal{V}_{p\geq a}$. Figure 13(a) shows the number of victims ($|\mathcal{V}_{p>a}|$) that present a collision probability greater than $a$, on the $x$ axis. Even at a small threshold $a=0.01$, we only find 664 accounts (0.04% of total victims). At first, excluding $\mathcal{V}_{p\geq a}$ does not seem to impact our detection performance. However, we find that excluding these accounts results in numerous false negatives because attackers purposefully target very active accounts, as shown in §6.1. Figure 13(b) illustrates the number of attacks we would report if we excluded attacks on $\mathcal{V}_{p>a}$. For example, we would discover 16.2M million poisoning transfers at $a=0.01$, which is 1 million fewer than at $a=0.99$. Indeed, we find that most of those detected attacks are indeed poisoning transfers. 99.7% of attacks targeting $\mathcal{V}_{p>0.99}$ are launched by the attack accounts $\mathcal{A}$ that have also targeted victims $\mathcal{V}_{p<0.01}$ (i.e., the population that is hardly susceptible to a collision).
False positives on payoff transfers are even more unlikely as they require all of the following: 1) the victim interacts with two similar addresses, 2) one of similar addresses “unintentionally” performs the poisoning transfer (e.g., testing with a zero transfer), and 3) the intended transfer, the poisoning transfer, and the payoff transfer happen in that correct order. As shown in §5.2, we do not find any false positives for top 30 cases for both chains.
The result shows that our system is robust against spurious collisions and that setting a low threshold reduces the performance of our detection system. Hence, we decide to set the threshold of $a=0.999$ to exclude victims that are interacting with a large enough number of addresses to be extremely susceptible to collisions, while capturing nearly all attacks. Setting such a high threshold only excludes, for instance, 64 Ethereum accounts.
(a) Nr. of accounts ($y$-axis) with a probability of collision greater than a given threshold ($x$-axis)
(b) Nr. of poisoning transfers detected at a given collision probability threshold
Figure 13: The number of victims that are susceptible to a collision and the number of attacks that are detected when we exclude those victims. The $x$-axis does not start with 0.
Appendix C Accidental transfers
@@
Measurement results. We identify 363 accidental transfers (i.e., not the result of a deliberate attack) from 295 victims and 336 mistyped destination addresses, totaling a 5.5M USD loss in Ethereum. This high number suggests that some victims have sent assets to the same mistyped address multiple times, or made typing mistakes multiple times, over different intended addresses. We observe a similar number of accidental transfers in BSC: 318 cases from 206 victims and 241 mistyped addresses for a total of 57,562 USD lost. By manually looking at accidental transfers, we confirmed that those transfers appear to be typing mistakes, and noticed common human errors. For instance, users 1) mistype a character with a neighboring key (i.e., low fat finger-distance [21]): “4” vs. “3,” “3” vs. “e,” 2) swap the order of two characters, or 3) mistype a character for something visually similar [32]: “b” vs “d,” “e” vs “c,” or “6” and “b.” To quantitatively assess typing mistakes, we calculated the Levenshtein distance $\ell$ (i.e., the number of operations necessary to replace a string with one another, including transposition) between two addresses. As expected, the distance is small; 84.95% of typos have $\ell=1$, and 96.69% have $\ell\leq 2$ in Ethereum. We have similar results in BSC (86.48% and 94.34% for $\ell=1$ and $\ell\leq 2$).
Appendix D Detecting copying bots
Here we explain why we use the attack ratio to exclude bots and show the robustness of our clustering results through temporal clustering.
As discussed in §4.2, we group two poisoning transfers if they 1) are in the same $TX$, 2) are launched by the same attacker $A$, or 3) specify the same lookalike address $L$. Our cluster considers two groups to be the same if they share even one link. We notice that two seemingly distinct large groups (based on their strategies and behaviors) are connected by a few attack addresses $A$ that behave differently from other attack addresses. We find that they appear to be the copying bots that duplicate transactions from multiple attack groups. For instance, the account 0x7D575a7C732D1c502c07f18BB822D29CF7DBf9E8 copies not only address poisoning transactions [^18] The original transaction: 0x725eaedf8857e243587020de97e-
@@ Appendix D Detecting copying bots
but also other transactions such as from MEV (Maximum Extractable Value) bots [^19] The original transaction: 0x0720fec6980c196b3545d273da2-
@@
. This type of copying behavior is relatively common in Ethereum or BSC, as discussed in previous work [27, 26].
Those bots seem to copy the original transactions without properly understanding address poisoning. Without changing the lookalike address $L$, the payoff transfer goes to the original attacker regardless of who initiates the attack. In other words, these bots are helping other attackers.
We next explain why we choose to use the attack ratio to detect copying bots over other methods (using the notion of time or removing all duplicated transactions).
Using the notion of time. Copy transactions generally come after the original poisoning transactions (in terms of the block number), so removing the second or later transactions could be effective. However, we notice that some copy transactions end up in the same block as the original transaction. This can be achieved by monitoring unconfirmed transactions in the peer-to-peer (P2P) network. For instance, the account 0xfffdcF2B3419C243A5eba5f051A64ad629362c9a manages to include copy transactions in the same block as the original transaction most of the time. Based on the public mempool data from Flashbots [^20] https://mempool-dumpster.flashbots.net/index.html, we realize that this account’s transactions do not appear in the public P2P network, while the original transactions do. This suggests that some bots monitor the original transactions in the P2P layer, and directly submit their transactions to the block builders (to potentially front-run the original transaction). As a result, we cannot systematically determine whether the first transaction is the original or the copy.
@@
We can also exclude duplicated transfers that have the same ($L$, $V$, $CT$ (or $AC$), and $v$ (value)). This method appears to capture all the transactions from the bots. However, we realize that the attackers (not the copying bots) also re-send their poisoning transfers several times against the same victim. We would remove more than 11 million (62%) “duplicated” transfers with many false positives, which significantly undermines the final clustering results (i.e., the size of the clusters becomes exceedingly small).
Removing potential bots by attack ratio. Only using information from transactions either leads to 1) the removal of too many non-bot transactions or 2) the inclusion of copying bot transactions. We decide to retrieve additional information of attack accounts $A$, and exclude $A$ with a low attack ratio—the ratio of the transactions used for address poisoning, as introduced in §4.2.
@@
Figure 14: Left: Top 10 clusters ($y$-axis) based on the threshold ($x$-axis). Right: the ratio of removed transfers over the total number of transfers based on the threshold
Finally, to additionally validate the robustness of our clustering, we perform a temporal clustering. We apply our clustering algorithm for the data until time $x$ and gradually increase $x$ over time to observe the change in the clustering results. Figure 15 shows how the top 10 clusters (based on the number of lookalike addresses) evolve over time where the link indicates the same attack group. The left figure shows the result with a threshold of 0.5; the clusters appear to be stable over time. The right figure is the result without any threshold (no bot exclusion); some large clusters remain separate for a while, but eventually connected around July and August 2023, forming a single large cluster due to the appearance of copying bots.
@@
Figure 15: Temporal clustering results. The x-axis is the time, and the y-axis is the number of clusters.
Appendix E Intended addresses
We focus here on the intended addresses $R$—the victim’s recipient which attackers try to impersonate. Table 7 lists the top ten intended addresses the attackers impersonate the most (based on the number of lookalike addresses $L$ targeting each $R$). We label each address based on the third-party Arkham Intel website. [^21] https://intel.arkm.com/ We also aggregate the number of attacks targeting each $R$. The table shows that many of $R$ are centralized exchange hot addresses. Since users do not directly interact with those hot wallets, we do not observe any successful cases for those addresses. Yet, those exchanges’ deposit addresses frequently interact with those hot wallets, which may explain why the attackers’ (algorithm) capture those intended addresses (as discussed in §6.2).
Table 7: Top 10 intended addresses
Intended address $R$ Label Num of $L$ Num of $TR$ 0xA9D1e0…1d3E43 Coinbase Hot Wallet 143,764 18,229 0x28C6c0…f21d60 Binance Hot Wallet 26,373 16,147 0x974CaA…ECC400 Stake.com Hot Wallet 3,254 2,355 0xF7C8dA…38d921 Devout Deposit 5,242 2,014 0xf89d7b…5EaA40 Bybit Hot Wallet 4,556 1,702 0xBFCd86…23d6aD HitBTC Hot Wallet 3,216 1,569 0x731309…AB4D2a BitGet Hot Wallet 7,547 1,477 0x99870D…9AF1e2 HitBTC Deposit 9,700 1,449 0x3CC936…aeCF18 MEXC Hot Wallet 6,603 1,397 0x75e89d…1dcB88 MEXC Hot Wallet 1,629 1,257
Appendix F UI Design of wallets and chain scanners

### v3

@@ 8 Mitigations
Protocol-level mitigations: One mitigation is to map human-readable strings to each address, like DNS for IP addresses. Some entities already have adopted domain names for wallet addresses through Blockchain Domain Name System (BNS). Human-readable addresses would reduce address poisoning as well as accidental transfers (Appendix D) and enhance usability. BNS, however, fosters other security risks such as typosquatting [22], name hijacking (i.e., re-using the expired domain names) [45], or name collisions [14].
@@
Contract-level mitigations: One solution is to disallow zero-value transfers in “transferFrom,” despite the existence of use cases (e.g., testing by developers). Specifically, we could theoretically modify the major stablecoin contracts (if they are upgradable) and mandate permission from the token sender, even when the transfer amount is zero. However, the community [^20] https://github.com/OpenZeppelin/openzeppelin-contracts/issues/3931 has concluded that no changes to these contracts should be made, as removing zero-value transfers (in transferFrom) would break existing code (e.g., “_becomeImplementation” in CDaiDelegate.sol (Compound) that relies on such zero-value transfers.)
Another potential (post-attack) solution is to quickly blacklist attack addresses. For the largest payoff transfer we observe, the Tether team (USDT) blocked the attacker’s address within an hour, which prevented the attacker from moving the retrieved funds to another address. [^21] Transaction blocking $L$: 0x2675aaf5db84f9f966c431ff82f2173d3251020b11b4939b4961ebedf92c78dd.
@@ 9 Related work
We highlight work particularly related to our discussion, with a focus on attack scale. Gao et al. [8] identified 2,117 counterfeit tokens by applying keyword matching to the 100 most popular tokens. The authors investigated two techniques involving counterfeit tokens (airdrop and arbitrage), leading to total financial losses exceeding 17M USD over 7,000 victims. Xia et al. [44] identified around 10,000 counterfeit tokens on Uniswap (50% of all Uniswap tokens) through ground truth labels provided by the platform, address clustering, and machine learning-based techniques, leading to a 16M USD loss from approximately 40,000 victims. He et al. [12] introduced TxPhishScope, a tool designed to detect blockchain phishing websites. The authors identified 26,333 phishing websites with 3,486 phishing blockchain accounts. Muzammil et al. [22] studied typosquatting on BNS, and identified over 26,000 typosquatting names and thousands of transactions sent to the addresses mapped to these squatting names. Other studies on free giveaway scams reported 24M–69.9M USD losses [19], including 872K USD losses from Twitter [17]. Based on these other works, blockchain address poisoning appears to be one of the largest (in terms of aggregate losses) and most widely targeted on-chain criminal activities.
Finally, we clarify our contributions in contrast to other contemporary works, most notably Guan and Li [10] and Ye et al. [47]. There are three major differences between our approach and Guan and Li’s. First, our clustering is more fine-grained. While we use guilt-by-association techniques to cluster attack instances (similar to theirs), crucially, we detect copying bots, which will erroneously merge distinct groups into one large cluster, as described in §4.2 and Appendix E. Second, we ran novel experiments and analyses, including victim characteristics, success conditions, group competitions, and address generation modeling, which provide deeper insights into attacker strategies and computational capability. Third, we uncovered 13 times more attacks (270M total) than prior works and compared the results between the two chains. Despite the similar chain structure, BSC requires more effort and optimization due to the large number of transfers. (Our BSC database is 8.6 times larger than Ethereum’s). That dataset also allows us to uncover cross-chain attacks.
Another contemporary piece, by Ye et al. [47], examines zero-value transfers, but not other types of poisoning transfers, which we do cover. In Appendix 7, we adjust our measurement period to provide a direct comparison of detection outcomes between our work and those two studies.
While most industry reports are case studies, a Chainalysis’s blog post [35] extended one case study (WBTC \$68M loss) to identify an additional 82,031 lookalike addresses. While commendable, this remains small compared to the whole ecosystem, as shown by our work.
@@ 11 Ethical considerations
While the concept of address poisoning is already the subject of many online discussions, we notified Etherscan and the MetaMask security team of our findings; MetaMask acknowledged that they are aware of address poisoning. Although attackers could potentially misuse some of our findings to improve their strategies, the main contribution of our paper is to demonstrate its practical severity in economic terms. We hope that our work can help users recognize when they are in targeted populations, and accordingly, take precautionary measures. We also introduce possible protection strategies. Our proposed mitigations can hopefully inspire wallet operators.
12 Open science
We released the detection result dataset, including over 17 million attack attempts on Ethereum and successful payoff transfers. [^22] https://kilthub.cmu.edu/articles/dataset/Blockchain_Address_Poisoning_Companion_Dataset_/29212703/1 We also provide a Jupyter notebook explaining 1) how to access the dataset, 2) how to produce descriptive statistics such as the number of poisoning transfers, and 3) how to manually verify the payoff transfer on Etherscan (BSCscan). This dataset will enable other researchers to validate our results as well as conduct further analysis.
@@ Acknowledgments
Appendix A Investigating false negatives
This appendix provides more details about the procedure we use to categorize false negatives in §5.2. We first obtain the set of victims $V$ that are seemingly attacked by lookalike addresses $L$ that we had missed in the Guan and Li’s dataset [10] we used for evaluation. We then retrieve each victim’s ERC-20 or native ETH transfers on Etherscan. We follow the flow chart laid out in Figure 13 and use the criteria defined in the green boxes to determine the reason (in the blue boxes) why we missed the lookalike address. “Mislabeling” indicates that the false negatives were actually true negatives – and that the issue lied with the dataset we compared to. (There are 45 such cases.) The asterisk (*) indicates that we lower the similarity threshold if we do not find the intended address $R$ with the current threshold $(3,4)$. The “unknown” category (orange box) signifies that the reason is (primarily) due to data collection quirks. For example, we fail to capture the original transfer between the victim $V$ and the intended address $R$ if the transfer is outside of our measurement period (2 years) or uses transferFrom instead of transfer. (We capture transferFrom when used for poisoning transfers.) These concern only 38 false negatives out of 164.
Figure 13: Flow chart of the identification of false negatives causes.
Appendix B Additional comparisons with other works
This appendix provides additional details on the performance comparison of our detection algorithm to Ye et al [47] and Guan and Li [10]. For a fair comparison, here, we restrict the data we collected to the same time period as those studies.
We detect more payoff transfers than Ye et al. We find 1,059 payoff transfers with 36.6M USD in losses while they found 560 transfers with at most 28M USD in losses, including two non-address poisoning attacks they investigate. For poisoning transfers, we identify 4.9M zero-value transfers, compared to 21M for Ye et al. It is unclear whether the difference comes from 1) false positives on Ye et al.’s side (i.e., benign zero-value transfers), or 2) false negatives on our side.
Table 7 summarizes our comparison with Guan and Li [10]. First, we find a similar total amount of payoff transfers, corroborating the validity of both approaches. Second, both approaches estimate similar attack costs in ETH, but not in USD. Guan and Li use a fixed ETH-USD conversion rate (around 3,200 USD) whereas we query the price at the time of attack. Because the exchange rate fluctuated between 1000 and 2000 USD for the majority of the observation period, Guan and Li overestimate attack costs by over 10 million USD, which results in a difference of a 300% on the return-on-investment (RoI). Third, we have more than 2.2 million counterfeit token transfers and 3,500 counterfeit tokens. Guan and Li appear to only report counterfeit tokens whose symbols are USDT or USDC. For example, out of 5,838 counterfeit tokens we find, only 1,922 use the exact same token symbols whereas at least 740 tokens use the different encodings of USDT or USDC (e.g., using a Cyrillic “S” and/or “T” in “USDT”). Finally, our technique misses attacks where $L$ matches $R$ in less than 7 digits. Guan and Li do consider these attacks, which, however, are shown to have a negligible impact in §5.2 (and which come at the increased risk of false positives).
@@ Appendix B Additional comparisons with other works
Payoff Cost ETH ROI Tiny Zero-value Counterfeit CT transfer (USD) token Guan and Li [10] 76.79M 7918.2 (25.5M) 2.2 277.3K 7.85M 6.64M 2,333 This paper 76.81M 8202 (14.9M) 5.1 283.9K 7.05M 8.87M 5,838
Appendix C Impact of the Birthday Paradox
@@
We examine whether our system should exclude a set of victims that present a collision probability greater than a given threshold $a$, i.e., $p\geq a$. We denote this set of victims by $\mathcal{V}_{p\geq a}$. Figure 14 shows the number of victims ($|\mathcal{V}_{p>a}|$) that present a collision probability greater than $a$, on the $x$-axis. Even at a small threshold $a=0.01$, we only find 664 accounts (0.04% of total victims). Excluding $\mathcal{V}_{p\geq a}$ does not seem to impact our detection performance. However, excluding these accounts results in numerous false negatives because attackers purposefully target very active accounts, as shown in §6.1. Figure 14 illustrates the number of attacks we would report if we excluded attacks on $\mathcal{V}_{p>a}$. For example, we would discover 16.2M million poisoning transfers at $a=0.01$, which is 1 million fewer than at $a=0.99$. Indeed, we find that most of those detected attacks are indeed poisoning transfers. 99.7% of attacks targeting $\mathcal{V}_{p>0.99}$ are launched by the same attack accounts $\mathcal{A}$ that have also targeted victims $\mathcal{V}_{p<0.01}$ (i.e., the population that is hardly susceptible to a collision).
False positives on payoff transfers are even more unlikely as they require: 1) the victim interacts with two similar addresses, 2) one of similar addresses unintentionally performs the benign transfer (e.g., a zero transfer for testing), and 3) the intended, the poisoning, and the payoff transfers happen in that correct order. We indeed do not find any false positives for the top 30 cases for both chains in §5.2.
Those results imply that our system is robust against spurious collisions and that setting a low threshold reduces the performance of our detection system. Hence, we decide to set $a=0.999$ to exclude victims that are interacting with a large enough number of addresses to be extremely susceptible to collisions, while capturing nearly all attacks. Setting such a high threshold only excludes 64 Ethereum accounts.
Figure 14: Left: Nr. of victims that are susceptible to a collision ($y$-axis) given a threshold ($x$-axis). Right: Nr. of attacks detected ($y$-axis) when we exclude those victims with a threshold ($x$-axis). The $x$-axis does not start with 0.
Appendix D Accidental transfers
@@
Measurement results. We identify 363 accidental transfers (i.e., not the result of a deliberate attack) from 295 victims and 336 mistyped destination addresses, totaling a 5.5M USD loss in Ethereum. This high number suggests that some victims have sent assets to the same mistyped address multiple times, or made typing mistakes multiple times, over different intended addresses. We observe a similar number of accidental transfers in BSC: 318 cases from 206 victims and 241 mistyped addresses for a total of 57,562 USD lost. By manually looking at accidental transfers, we confirmed that those transfers appear to be typing mistakes, and noticed common human errors. For instance, users 1) mistype a character with a neighboring key (i.e., low fat finger-distance [21]): “4” vs. “3,” “3” vs. “e,” 2) swap the order of two characters, or 3) mistype a character for something visually similar [33]: “b” vs “d,” “e” vs “c,” or “6” and “b.” To quantitatively assess typing mistakes, we calculate the Levenshtein distance $\ell$ (i.e., the number of operations necessary to replace a string with one another, including transposition) between two addresses. As expected, the distance is small; 84.95% of typos have $\ell=1$, and 96.69% have $\ell\leq 2$ in Ethereum. We have similar results in BSC (86.48% and 94.34% for $\ell=1$ and $\ell\leq 2$).
Appendix E Detecting copying bots
Here we explain why we use the attack ratio to exclude bots and show robustness in our results via temporal clustering.
As discussed in §4.2, we group two poisoning transfers if they 1) are in the same $TX$, 2) are launched by the same attacker $A$, or 3) specify the same lookalike address $L$. Our cluster considers two groups to be the same if they share even one link. We notice that two seemingly distinct large groups (based on their strategies and behaviors) are connected by a few attack addresses $A$ that behave differently from other attack addresses. We find that they appear to be the bots that simply copy transactions from multiple attack groups. For instance, the account 0x7D575a7C732D1c502c07f18BB822D29CF7DBf9E8 copies not only address poisoning transactions [^24] The original transaction: 0x725eaedf8857e243587020de97e-
@@ Appendix E Detecting copying bots
but also other transactions such as from MEV (Maximum Extractable Value) bots. [^25] The original transaction: 0x0720fec6980c196b3545d273da2-
@@
This type of copying behavior is relatively common in Ethereum or BSC, as discussed in previous work.[28, 27]
Those bots seem to copy the original transactions without properly understanding address poisoning. Without changing the lookalike address $L$, the payoff transfer goes to the original attacker regardless of who initiates the attack. These bots are helping other attackers. We next explain why we use the attack ratio to detect copying bots over other methods (using the notion of time or removing all duplicated transactions).
Using the notion of time. Copy transactions generally come after the original poisoning transactions (in terms of the block number), so removing the second or later transactions could be effective. However, we notice that some copy transactions end up in the same block as the original transaction. This can be achieved by monitoring unconfirmed transactions in the peer-to-peer (P2P) network. For instance, the account 0xfffdcF2B3419C243A5eba5f051A64ad629362c9a manages to include copy transactions in the same block as the original transaction. Based on the public mempool data from Flashbots, [^26] https://mempool-dumpster.flashbots.net/index.html we realize that this account’s transactions do not appear in the public P2P network, while the original transactions do. This implies that some bots monitor the original transactions in the P2P layer, and directly submit their transactions to the block builders (to front-run the original transaction). As a result, we cannot systematically determine whether the first transaction is the original or the copy.
@@
We can also exclude duplicated transfers that have the same ($L$, $V$, $CT$ (or $AC$), and $v$ (value)). This method appears to capture all the transactions from the bots. However, we realize that attackers (not the copying bots) also resend their poisoning transfers several times against the same victim. We would remove more than 11 million (62%) “duplicated” transfers, which significantly undermines the final clustering results (i.e., the size of the clusters becomes exceedingly small).
Removing potential bots by attack ratio. Only using attack transactions either leads to 1) the removal of too many non-bot transactions or 2) the inclusion of copying bot transactions. We decide to retrieve additional information of attack accounts $A$, and exclude $A$ with a low attack ratio—the ratio used for address poisoning, as introduced in §4.2.
@@
Figure 15: Left: Top 10 clusters ($y$-axis) based on the threshold ($x$-axis). Right: the ratio of removed transfers over the total number of transfers based on the threshold.
To additionally validate our clustering robustness, we perform a temporal clustering. We apply our algorithm for the data until time $x$ and gradually increase $x$ over time to observe the change in the results. Figure 16 shows how the top 10 clusters (based on the number of lookalike addresses) evolve over time where the link indicates the same attack group. The red line denotes Group 1; the blue lines represent other groups. The left figure shows results with an attack ratio threshold of 0.5; the clusters appear to be stable over time. The right figure shows results without any threshold (i.e., we do not exclude bots); large clusters remain separate for a while, but eventually get connected around Jul. and Aug. 2023, forming a single large cluster due to the emergence of copying bots.
@@
Figure 16: Temporal clustering results. $x$-axis is time, $y$-axis is cluster rank of at time $x$. Red: Group 1; blue: other groups. Links are used to denote the same group.
Appendix F Intended addresses
In this section, we focus on the intended addresses $R$—the victim’s recipient, which attackers try to impersonate. Table 8 lists the top ten intended addresses that the attackers impersonate the most (based on the number of lookalike addresses $L$ targeting each $R$). We label each address based on the third-party Arkham Intel website, [^27] https://intel.arkm.com/ and aggregate the number of attacks targeting each $R$. Table 8 shows that many of $R$ are centralized exchange hot addresses. Since users do not directly interact with those hot wallets, we do not observe any successful cases for those addresses. Yet, those exchanges’ deposit addresses frequently interact with those hot wallets, which may explain why the attackers’ (algorithm) capture those intended addresses (as discussed in §6.2).
Table 8: Top 10 most targeted intended addresses
Intended address $R$ Label Num of $L$ Num of $TR$ 0xA9D1e0...1d3E43 Coinbase Hot Wallet 143,764 18,229 0x28C6c0...f21d60 Binance Hot Wallet 26,373 16,147 0x974CaA...ECC400 Stake.com Hot Wallet 3,254 2,355 0xF7C8dA...38d921 Devout Deposit 5,242 2,014 0xf89d7b...5EaA40 Bybit Hot Wallet 4,556 1,702 0xBFCd86...23d6aD HitBTC Hot Wallet 3,216 1,569 0x731309...AB4D2a BitGet Hot Wallet 7,547 1,477 0x99870D...9AF1e2 HitBTC Deposit 9,700 1,449 0x3CC936...aeCF18 MEXC Hot Wallet 6,603 1,397 0x75e89d...1dcB88 MEXC Hot Wallet 1,629 1,257
Appendix G UI Design of wallets and chain scanners